### v1

@@ 2.1 Problem Statement
An Elastic Rod can be kinematically described through a time-variant 3D curve, usually called backbone. The length of this curve can be parameterized by a material curvilinear abscissa . Every point of the backbone can be written w.r.t. an inertial frame . Usually, in literature, the point of the backbone at is called base and at is called tip of the robot. Furthermore, every cross-section has associated geometrical and material quantities: the cross-section area , the second moment of area about the local axes ; the mass density of the rod , the Young Modulus of the material , and the shear Modulus , where is the Poisson coefficient.
@@ 2.2 Cosserat Rod Theory
The Cosserat rod theory (CRT) (Cosserat and Cosserat, 1896) describes all strain modes of the rod. This model is useful for describing CSM that interact with cluttered environments, in which shear and elongation deformations play a fundamental role. Each cross-section is associated with a reference system , whose axes are called directors. The relative roto-translation between and is expressed by the homogeneous matrix defined as
@@ 2.2.1 Kinematics
Definition (5) indicates the infinite DoFs of an elastic rod. The Configuration Space results in a functional space of curves in (Boyer et al., 2020). A Configuration Space with an infinite dimension is a unique feature of soft robots, making control very challenging.
@@ 2.2.2 Dynamics
After defining these quantities, it is possible to compute the Dynamics of an elastic rod using the Poincaré equations (Renda et al., 2018a)
The EoM (8) is a set of partial differential equations of a Cosserat Rod in the local frame. It is common also to find the same set of PDE in an explicit form and expressed in the inertial frame, such as
@@ 2.2.3 Strain Parameterization
The Configuration Space (5) and PDE (8) show the infinite DoFs of an elastic rod. From a control and numerical implementation perspective, it is convenient to find a method to discretize the continuous nature of the soft segment. Renda et al. (2020) and Boyer et al. (2020) proposed a Strain Parameterization, expressing the Kinematics and the Dynamics of the elastic rod in terms of strain twist . Let us again consider the definition of the strain twist (2), which is easily rewritten as . Assuming that the initial condition is known, the function can be univocally determined by the strain twist. The Configuration Space can be considered the Shape Space of the elastic rod, which is a functional space of the -parameterized curves in , such as
@@
where and is a vector of generalized coordinates, with . The main idea of the discretization technique is to truncate the basis matrix to a finite number of columns, reducing the discretized Shape Space to . This approach allows the user to choose the degree of approximation or neglect certain strain modes. To solve the PDE, the proposed method uses the Magnus Expansion (Haier et al., 2006), resulting in a convenient Product of Exponentials, which is widely used in classical robotics (Murray et al., 1994)
@@
The EoM (14) can be solved using a standard time solver, such as Runge-Kutta or Explicit Euler, as implemented in the SoRoSim simulator (Mathew et al., 2022).
@@ 2.2.4 Discrete Elastic Rod
Another method to discretize the continuum nature of elastic rods was introduced in the pioneering work of Gazzola et al. (2018). First, the Authors derived the EoM (9), including an elongation/compression ratio defined as , where is the curvilinear abscissa in the rest configuration. This ratio is present because the length is parameterized using the curvilinear abscissa differently from (9). In the presence of axial stretching, the stretching ratio scales the geometrical quantities
@@
It is worth highlighting that (16) contains an additional contribution with respect to (8), which depends on the time derivative of stretching ratio . Furthermore, scaling the geometrical quantities partially relaxes the assumption of rigid cross-sections. To numerically resolve the EoM (16), they extended for a Cosserat rod the spatial discretization algorithm proposed in a previous study (Bergou et al., 2008), which discretizes the rod in a sequence of rigid segments connecting nodes. For each node, the EoM consider the interactions with the other nodes and external forces. In addition, it is possible to associate kinematic and dynamic quantities to each node and segment to solve (16). In the discrete domain, some quantities, such as curvature, must be expressed in an integrated form over the domain (Gazzola et al., 2018). This domain corresponds to the Voronoi region associated with the interior nodes ,
@@
where is the point-wise mass associated with the node, is a domain dilation factor, and are the discrete difference and the averaging operator defined in (Gazzola et al., 2018). In the first equation of (19), , and . In contrast, in the second equation, ranges , while still ranges from . This distinction arises from the definition of discrete curvature and bending stiffness within the interior nodes set.
For the time integration of the discretized EoM (19), the authors proposed the use of a Second-Order Position Verlet time integrator, which exhibits a good balance between numerical accuracy and computational cost (Gazzola et al., 2018). To ensure numerical stability, they proposed an empirical law to choose the integration frequency as . The discrete EoM (19) were implemented in the PyElastica simulator (Naughton et al., 2021).
@@ 2.3 Kirchhoff Rod Theory
The Kirchhoff rod theory (KRT) (Love, 1906) is a special case of CRT that considers an elastic rod unstretchable and unshearable. It is particularly suitable for CSM that bend around any axis and twist. Notably, it introduces the notion of a directed curve, assigning a specific reference system to each cross-section. The EoM of Kirchhoff rods can be computed by specializing the EoM of Cosserat, that is (8) or (9). In particular, the constraint of an unstretchable and unshearable rod can be written as
This constraint can be added in the EoM (16), resulting in
The linear internal force serves as a Lagrangian multiplier and it is a virtual internal force that constrains the rod from stretching or shearing. In the case of Strain Parameterization, it is sufficient to apply the constraint (20) in the strain twist
Therefore, the basis function matrix only generates the angular strain mode vector .
@@ 2.4 Euler-Bernoulli Rod Theory
The Euler-Bernoulli rod theory (EBRT) (Timoshenko, 1983) is one of the simplest rod theories in which the rod can only bend around one axis. It can be considered the 2D case of the KRT without the twist. Here, the assumption is that the slope angle of the rod is equal to the tangent angle of the backbone curve. Below, we report linear and nonlinear versions.
@@ 2.4.1 Linear Euler-Bernoulli
where is the -component of the internal moment . For control purposes, the linearity of (24) facilitates the use of efficient controllers from the classic control theory (Doroudchi et al., 2018).
@@ 2.4.2 Nonlinear Euler-Bernoulli
In the nonlinear EBRT (or Euler’s Elastica), the rod length is parameterized by the arclength . Recalling the Fundamental Theorem of the local theory of Curves in the 2D case, every regular curve can be determined by the curvature (do Carmo, 1976). From this concept, Della Santina and Rus (2019) reformulated Elastica in a classical robotic formulation. Let be the -component of the curvature twist . The Cartesian pose of every cross-section is uniquely determined by the curvature, i.e.
@@
The geometrical meaning of the polynomial curvature is to constrain the shape of the backbone curve to be a Generalized Cornu Spiral, i.e., a curve with a polynomial curvature (Della Santina and Rus, 2019). It is worth highlighting that the (27) is equivalent to write (11) with a polynomial basis function, related only to the curvature (Caradonna et al., 2024). Defining as joint variables, it is possible to write the EoM as
@@ 2.5 Timoshenko Rod Theory
The Timoshenko-Ehrenfest rod theory (Timoshenko and Goodier, 1951) extends the EBRT, relaxing the unshearability constraint. Therefore, the equality between the tangent angle and the cross-sections’ angle is no longer valid (i.e., ). For this reason, the EoM of a Timoshenko rod considers the angle as an independent variable. Its dynamics can be written as
@@ 3.1 The Actuation Matrix
Consider a Cosserat rod with actuators. Let be the distance from the center of the cross-section and the -th actuator (Fig. 3). Each actuator applies an internal active wrench that depends on the actuator routing. We can express the internal active wrench applied by the actuator as
where is the tangent versor to the actuator path and is the magnitude of the i-th actuator. For example , can be
where is the cable tension, is the pneumatic pressure and is the internal cross-section area of the fluidic chamber. Moreover, the expression of the tangent versor is derived as
@@ 3.2 Actuators-Strain Mapping
From the definition of actuation matrix (35), it is possible to map the actuation magnitude vector to the distributed active load . Renda et al. (2020) propose a statics-based method to relate the actuators with excited strain modes. In particular, let be the statics of a Cosserat Rod as
@@
From this equality, Renda et al. (2024) propose a trivial form of the static equation, choosing the implicit parametrization
@@
The functional basis derived from the implicit parametrization (39) uses the information of actuation routing contained in , to provide the minimum set of functional basis. In particular, these functional bases describe which are the excited strain modes and which is the best function to describe them statically. Furthermore, in (35) considers also the geometrical and the material information of the robot, contained in . Finally, (39) can be a useful starting point to build the optimal functional basis matrix , avoiding redundant or inefficient shape functions, preserving accuracy and computational efficiency.
@@ 4 Rod Models for Continuum and Soft Robots
The rod models were classified based on combinations of the four deformation modes. Out of the fifteen possible combinations, only nine classes emerged. Fig. 4 shows the distribution of rod models over deformations classes. Within each deformation class, we divided the papers mainly by actuation, robot type, or topic. In each division, the articles follow a chronological or logical order. An advanced Scopus query aided the selection of the articles (see Appendix A). Fig. 5 shows how different robots might exhibit these deformations induced by actuators or the interaction with the environment. We survey rod models ordering the deformation classes from most to least popular (see Fig. 4).
We summarize the four major deformation classes (i.e., Bend, All Deformation, Bend & Twist, Bend & Stretch) respectively in Table 1-4 to facilitate the literature comprehension. Each table arbitrarily reports representative articles to distill the deformation class and highlight the versatility of rod models.

### v2

@@ 4.1 Problem Statement
An Elastic Rod can be kinematically described through a time-variant curve, usually called backbone. The length of this curve can be parameterized by a material curvilinear abscissa . Every point of the backbone can be written w.r.t. an inertial frame . Usually, in literature, the point of the backbone at is called base and at is called tip of the robot. Every cross-section has associated geometrical and material quantities: the cross-section area , the second moment of area about the local axes , the mass density of the rod , the Young Modulus of the material , and the shear Modulus , where is the Poisson coefficient.
@@ 4.2 Cosserat Rod Theory
The CRT (cosserat1896) describes all strain modes of the rod. This model is useful for describing CSM that interact with cluttered environments, in which shear and elongation deformations play a fundamental role. Each cross-section is associated with a reference system , whose axes are called directors. The relative roto-translation between and is expressed by the homogeneous matrix defined as
@@ 4.2.1 Kinematics
Definition (5) indicates the infinite DoFs of an elastic rod. The Configuration Space results in a functional space of curves in (boyer2020dynamics). A Configuration Space with an infinite dimension is a unique feature of soft robots, making control very challenging.
@@ 4.2.2 Dynamics
For clarity, the distributed wrenches (, , ) represent the forces and torques acting on a particular cross-section (), rather than force or torque densities distributed along the body’s length.
After defining these quantities, it is possible to compute the Dynamics of an elastic rod using the Poincaré equations (renda2018discrete)
The EoM (8) is a set of PDE of a Cosserat Rod in the local frame. It is common also to find the same set of PDE in an explicit form and expressed in the inertial frame, such as (gazzola2018forward)
@@ 4.2.3 Strain Parameterization
The Configuration Space (5) and PDE (8) show the infinite DoFs of an elastic rod. From a control and numerical implementation perspective, it is convenient to find a method to discretize the continuous nature of the soft segment. renda2020geometric and boyer2020dynamics proposed a Strain Parameterization, also known as Geometrical Variable Strain (GVS) mathew2025reduced, expressing the Kinematics and the Dynamics of the elastic rod in terms of strain twist . Let us again consider the definition of the strain twist (2), which is easily rewritten as . Assuming that the initial condition is known, the function can be univocally determined by the strain twist. The Configuration Space can be considered the Shape Space of the elastic rod, which is a functional space of the -parameterized curves in , such as
@@
where and is a vector of generalized coordinates, with . The main idea of the discretization technique is to truncate the basis matrix to a finite number of columns, reducing the discretized Shape Space to . This approach allows the user to choose the degree of approximation or neglect certain strain modes. To solve the PDE, the proposed method uses the Magnus Expansion (harier2006geometric), resulting in a convenient Product of Exponentials, which is widely used in classical robotics (murray)
@@
The EoM (14) can be solved using a standard time solver, such as Runge-Kutta or Explicit Euler, as implemented in the SoRoSim simulator (mathew2022sorosim).
Recently, the Authors showed how to compute the analytical derivatives of the GVS approach mathew2024analytical, evolving SoRoSim as a differentiable simulator. In this work, the Authors also showed the benefits of the implicit time-integration schemes, such as the Implicit Euler and the Newmark- methods.
Finally, to fully describe soft robots interacting with the environment, xun2024cosserat proposed an extension of (14), including frictional contacts with rigid and soft bodies.
@@ 4.2.4 Discrete Elastic Rod
Another method to discretize the continuum nature of elastic rods was introduced in the pioneering work of gazzola2018forward. First, the Authors derived the EoM (9), including an elongation/compression ratio defined as , where is the curvilinear abscissa in the rest configuration. This ratio is present because the length is parameterized using the curvilinear abscissa differently from (9). In the presence of axial stretching, the stretching ratio scales the geometrical quantities
@@
It is worth highlighting that (16) contains an additional contribution with respect to (8), which depends on the time derivative of stretching ratio . Furthermore, scaling the geometrical quantities partially relaxes the assumption of rigid cross-sections. To numerically resolve the EoM (16), they extended for a Cosserat rod the spatial discretization algorithm proposed in a previous study (bergou2008discrete), which discretizes the rod in a sequence of rigid segments connecting nodes. For each node, the EoM consider the interactions with the other nodes and external forces. In addition, it is possible to associate kinematic and dynamic quantities to each node and segment to solve (16). In the discrete domain, some quantities, such as curvature, must be expressed in an integrated form over the domain (gazzola2018forward). This domain corresponds to the Voronoi region associated with the interior nodes ,
@@
where is the point-wise mass associated with the node, is a domain dilation factor, and are the discrete difference and the averaging operator defined in (gazzola2018forward). In the first equation of (19), , and . In contrast, in the second equation, ranges , while still ranges from . This distinction arises from the definition of discrete curvature and bending stiffness within the interior nodes set.
For the time integration of the discretized EoM (19), the authors proposed the use of a Second-Order Position Verlet time integrator, which exhibits a good balance between numerical accuracy and computational cost (gazzola2018forward). To ensure numerical stability, instead of enforcing rigorous Courant-Friedrichs-Levy stability conditions (courant1967partial), they proposed an empirical law to choose the integration time step as . The discrete EoM (19) were implemented in the PyElastica simulator (naughton2021elastica).
@@ 4.3 Kirchhoff Rod Theory
The Kirchhoff rod theory (KRT) (Love1906) is a special case of CRT that considers an elastic rod unstretchable and unshearable o2017modeling; gazzola2018forward; coleman1993dynamics. It is particularly suitable for CSM that bend around any axis and twist. Notably, it introduces the notion of a directed curve, assigning a specific reference system to each cross-section. The EoM of Kirchhoff rods can be computed by specializing the EoM of Cosserat, that is (8) or (9). In particular, the constraint of an unstretchable and unshearable rod can be written as
The constraint (20) enforces , which implies (gazzola2018forward). Incorporating this constraint into the EoM leads to the following modified equations of motion:
The linear internal force serves as a Lagrangian multiplier and it is a virtual internal force that constrains the rod from stretching or shearing. As stated in gazzola2018forward, inserting the constraint of an unshearable and unstretchable rod can increase the computational time.
In the case of Strain Parameterization, it is sufficient to apply the constraint (20) in the strain twist
Therefore, the basis function matrix only generates the angular strain mode vector . Unlike the previous approach, (22) does not apply as a constraint, and the EoM can be solved without considering any Lagrangian multipliers.
@@ 4.4 Euler-Bernoulli Rod Theory
The Euler-Bernoulli rod theory (EBRT) (timoshenko1983history) is one of the simplest rod theories in which the rod can only bend around one axis. It can be considered the 2D case of the KRT without the twist. Here, the assumption is that the slope angle of the rod is equal to the tangent angle of the backbone curve. Below, we report linear and nonlinear versions.
@@ 4.4.1 Linear Euler-Bernoulli
where is the -component of the internal moment . For control purposes, the linearity of (24) facilitates the use of efficient controllers from the classic control theory (doroudchi2018decentralized).
@@ 4.4.2 Nonlinear Euler-Bernoulli
In the nonlinear EBRT (or Euler’s Elastica), the rod length is parameterized by the arclength . Recalling the Fundamental Theorem of the local theory of Curves in the 2D case, every regular curve can be determined by the curvature (carmo_differential_geometry). From this concept, della2019control reformulated Elastica in a classical robotic formulation. Let be the -component of the curvature twist . The Cartesian pose of every cross-section is uniquely determined by the curvature, i.e.
@@
The geometrical meaning of the polynomial curvature is to constrain the shape of the backbone curve to be a Generalized Cornu Spiral, i.e., a curve with a polynomial curvature (della2019control). It is worth highlighting that the (27) is equivalent to write (11) with a polynomial basis function, related only to the curvature (caradonna2024model). Defining as joint variables, it is possible to write the EoM as
@@ 4.5 Timoshenko Rod Theory
The Timoshenko-Ehrenfest rod theory (Timoshenko1951) extends the EBRT, relaxing the unshearability constraint. Therefore, the equality between the tangent angle and the cross-sections’ angle is no longer valid (i.e., ). For this reason, the EoM of a Timoshenko rod considers the angle as an independent variable. Its dynamics can be written as
@@ 5.1 The Actuation Matrix
Consider a Cosserat rod with actuators. Let be the distance from the center of the cross-section and the -th actuator (Figure 3). Each actuator applies an internal active wrench that depends on the actuator routing. We can express the internal active wrench applied by the actuator as
where is the tangent unit vector to the actuator path and is the magnitude of the i-th actuator. For example, can be
where is the cable tension, is the pneumatic pressure and is the internal cross-section area of the fluidic chamber. The expression of the tangent unit vector is derived as
@@ 5.2 Actuators-Strain Mapping
From the definition of actuation matrix (35), it is possible to map the actuation magnitude vector to the distributed active load . renda2020geometric propose a statics-based method to relate the actuators with excited strain modes. In particular, let be the statics of a Cosserat Rod as
@@
From this equality, renda2024dynamics propose a trivial form of the static equation, choosing the implicit parametrization
@@
The functional basis derived from the implicit parametrization (39) uses the information of actuation routing contained in , to provide the minimum set of functional basis. In particular, these functional bases describe which are the excited strain modes and which is the best function to describe them statically. Furthermore, in (35) considers also the geometrical and the material information of the robot, contained in . Finally, (39) provides a useful starting point for constructing the optimal functional basis matrix , avoiding redundant or inefficient shape functions while preserving accuracy and computational efficiency.
@@ 6 Rod Models for Continuum and Soft Robots
We categorize rod models based on the deformation modes they represent, independently of the underlying actuation technology. From a control perspective, the specific actuation method (e.g., tendons, pneumatic chambers) is secondary to the type of strain it produces, such as elongation or contraction, torsion, or bending. The admissible deformation modes also implicitly determine the suitable rod theory. For example, a robot limited to planar bending is consistent with the Euler-Bernoulli framework, while more complex deformation behaviors might require Cosserat theory. While several studies formally employ the Cosserat rod formulation, in practice, they restrict the model to a limited subset of strain modes, effectively reducing it to simpler formulations. By explicitly classifying models according to the combination of strain modes they allow, we highlight such modeling choices. Out of the fifteen possible combinations of deformation modes, only nine classes emerged. We summarize the deformation classes in Table 2 to facilitate the literature comprehension, where we arbitrarily report representative articles to distill each class and highlight the versatility of rod models. An advanced Scopus query aided the initial selection of the articles (see Appendix A). We also discussed articles beyond the query.